# Dataset Condensation for Time Series Classification via Dual Domain Matching

### v1

@@ Abstract.
Dataset Condensation, Data Mining, Time Series Classification
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; June 2 2023; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data mining
@@ 1. Introduction
The exponential growth of time series data across various domains, such as traffic forecasting (Li et al., 2017; Liu et al., 2023a, b, 2024), clinical diagnosis (Hayat et al., 2022; Harutyunyan et al., 2019), and financial analysis (Cheng et al., 2022; Kumbure et al., 2022), has presented opportunities for researchers and practitioners. However, the sheer volume of data has imposed burdens in terms of storage, processing, and analysis in terms of deep learning context. In the context of some deep learning downstream tasks such as neural architecture search (Rakhshani et al., 2020) and continual learning (Gupta et al., 2021), the utilization of the full dataset for training has the potential to yield bad efficiency. Consequently, there is an urgent need to develop effective strategies for condensing time series datasets while preserving their essential information to alleviate the efficiency challenge of training a deep model.
@@ 17      Back-propagate $\mathcal{L}$ and update $\mathcal{S}$
5. Experiment
Table 2. Details of time series datasets.
Dataset # Train # Test Length # Channel # Class
HAR 7,352 2,947 128 9 6
Electric 8,926 7,711 96 1 7
Insect 25,000 25,000 600 1 10
FD 8,184 2,728 5,120 1 3
Sleep 25,612 8,910 3,000 1 5
@@ 5.2. Experiment Setting
Hyper-parameters: There are only a small number of hyper-parameters associated with CondTSC. For simplicity, we use the same set of hyper-parameters for all datasets. For the training process in Train Module, we use a learnable learning rate of 1e-3 following previous settings (Cazenavette et al., 2022; Zhang et al., 2023; Du et al., 2023). Moreover, we set $N=10$ and $M=1000$ for the number of the training iterations for $\mathcal{S}$ and $\mathcal{\mathcal{T}}$ respectively. For the loss in Dual Matching Module, we set $\lambda=1.0$. For the update of the synthetic data $\mathcal{S}$, we use an SGD optimizer with a learning rate of $lr_{\mathcal{S}}=1.0$ and update epochs $Epo=2000$. The whole experiment is implemented by Pytorch 1.10.0 on RTX3090. The code and data are available in https://github.com/TimeSeriesCond/TimeSeriesCondensation.

### v3

@@ Abstract.
Dataset Condensation; Data Mining; Time Series Classification
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671675isbn: 979-8-4007-0490-1/24/08ccs: Information systems Data mining
@@ 1. Introduction
The exponential growth of time series data across various domains, such as traffic forecasting (Li et al., 2017; Liu et al., 2023a, b, 2024a, 2024b; Liang et al., 2023), clinical diagnosis (Hayat et al., 2022; Harutyunyan et al., 2019), and financial analysis (Cheng et al., 2022; Kumbure et al., 2022), has presented opportunities for researchers and practitioners. However, the sheer volume of data has imposed burdens in terms of storage, processing, and analysis in terms of deep learning context. In the context of some deep learning downstream tasks such as neural architecture search (Rakhshani et al., 2020) and continual learning (Gupta et al., 2021), the utilization of the full dataset for training has the potential to yield bad efficiency. Consequently, there is an urgent need to develop effective strategies for condensing time series datasets while preserving their essential information to alleviate the efficiency challenge of training a deep model.
@@ 18      Back-propagate $\mathcal{L}$ and update $\mathcal{S}$
5. Experiment
Table 1. Details of time series datasets.
Dataset # Train # Test Length # Channel # Class
HAR 7,352 2,947 128 9 6
Electric 8,926 7,711 96 1 7
Insect 25,000 25,000 600 1 10
FD 8,184 2,728 5,120 1 3
Sleep 25,612 8,910 3,000 1 5
@@ 5.2. Experiment Setting
Hyper-parameters: There are only a small number of hyper-parameters associated with CondTSC. For simplicity, we use the same set of hyper-parameters for all datasets. For the training process in Train Module, we use a learnable learning rate of 1e-3 following previous settings (Cazenavette et al., 2022; Zhang et al., 2023; Du et al., 2023). Moreover, we set $N=10$ and $M=1000$ for the number of the training iterations for $\mathcal{S}$ and $\mathcal{\mathcal{T}}$ respectively. For the loss in Dual Matching Module, we set $\lambda=1.0$. For the update of the synthetic data $\mathcal{S}$, we use an SGD optimizer with a learning rate of $lr_{\mathcal{S}}=1.0$ and update epochs $Epo=2000$. The whole experiment is implemented by Pytorch 1.10.0 on RTX3090. The code and data are available in https://github.com/zhyliu00/TimeSeriesCond.
@@ 6. Conclusion
Acknowledgment
We wish to express our sincere appreciation to the researchers of (Dau et al., 2018) for their exceptional efforts in creating, curating, and maintaining this invaluable collection of time series datasets. This work was sponsored by National Key Research and Development Program of China under Grant No.2022YFB3904204, National Natural Science Foundation of China under Grant No. 62102246, 62272301, 62176243, and Provincial Key Research and Development Program of Zhejiang under Grant No. 2021C01034. Part of the work was done when the students were doing internships at Yunqi Academy of Engineering.